John and Joan have been Married for exactly 37 years. Joan is four years older than John. Now the sum of their age is 126. How old was Joan when they were married?
A.) Set up and write an equation that represents the unknown ages in terms of Joan’s age when they were married. ( The variable should represent Joan’s age when they were married)
B.) Determine how old Joan was when Joan and John were married.

Found 2 solutions by richwmiller, ikleyn:
Answer by richwmiller(17219)

a=Joan
h=John
m=a-37
h=a-4
a+h=126
solve for m
You need to combine all these equations into one equation for part A
2a-4-126
2a=130
a=65 Joan is 65 now
m=65-37
m=28 Joan was 28 when she married John
h=61 John is 61 now




Answer by ikleyn(52781)

.

            The given instructions in the post only make the solution harder.

            Let me to present you my solution in much more simple way.


Let x be the John age.

Then the Joan age is (x+4) years, according the condition.


We are given  

    x + (x+4) = 126.

Hence,  

    2x + 4 = 126,

    2x     = 126 - 4 = 122,

     x               = 122/2 = 61.


37 years ago, when they were married,  John was  61 - 37 = 24 years old.

Joan was  24+4 = 28 years old at that time.    ANSWER

-----------------

It is a bad teacher who instructs to solve a problem in a hard way, without any reason.